# Taming Modality Entanglement in Continual Audio-Visual Segmentation

### v1

@@ III-D Collision-based Sample Rehearsal
LwF [2] 10.1 25.1 12.3 7.1 19.0 8.8 1.5 9.7 2.6 1.5 12.6 3.0 1.3 0.7 1.3 1.3 4.5 1.6
LwF-MC [35] 2.0 2.2 2.0 16.4 1.1 14.3 2.8 0.03 2.4 5.8 0.6 5.0 1.6 0.0 1.5 1.3 1.7 1.4
ILT [36] 12.3 19.7 13.4 14.5 13.8 14.4 8.6 7.2 8.4 2.0 11.4 3.4 1.3 0.6 1.2 1.3 3.7 1.5
MiB [5] 17.4 23.0 18.2 17.5 16.6 17.4 4.1 11.5 5.1 5.7 7.3 5.9 1.6 2.8 1.7 1.3 4.9 1.5
PLOP [18] 21.2 13.5 20.1 19.0 11.3 17.9 1.3 11.7 10.0 8.3 9.3 8.4 1.3 0.2 1.2 1.2 4.1 1.4
AVSegFormer [37] 1.5 34.6 6.1 1.5 22.7 4.5 1.4 34.9 4.0 1.5 9.1 2.5 1.3 0.3 1.3 1.3 3.7 1.5
EIR [22] 14.6 1.3 12.8 12.4 0.1 10.7 6.8 1.1 6.0 5.5 0.2 4.8 0.5 0.08 0.4 0.5 0.02 0.4
@@
IV Experiments
IV-A AVSBench Datasets
In our work, a class-incremental audio-visual segmentation dataset (AVSBench-CI) is constructed from the well-known dataset AVSBench-semantic [10] to validate the proposed CMR. AVSBench-semantic utilizes the techniques introduced in VGGSound [38] to collect videos, ensuring that the audio and visual clips align with the intended semantics. The dataset provides semantic segmentation maps for videos as labels to enhance audio-visual semantic segmentation (AVSS). It contains a total of 11,356 videos spanning 70 categories. Each video segment consists of 10 frames of images and one 10-second audio clip. We divide the 70 categories in AVSBench-semantic for the original dataset into three training steps: 60-10, 60-5, and 65-1. Following the conventional continual semantic segmentation setup, the three training steps are divided into overlapped and disjoint settings to evaluate the model’s performance under different task stream configurations.
In the overlapped setting, the classes are divided sequentially, meaning that classes from past and future tasks may appear in the current data and be labelled as background. In the disjoint setting, a community detection algorithm [39] is employed to minimize the overlap of training data between consecutive steps. Therefore, the current data will not contain classes from future or past tasks. This setup closely aligns with continual learning scenarios. Furthermore, we expand the single-semantic dataset (AVSBench-CIS) and the multi-semantic dataset (AVSBench-CIM) based on the number of targets in the videos. AVSBench-CIS and AVSBench-CIM address scenarios involving modality entanglement with single-target and multi-target settings, respectively. The same settings are applied to these datasets.
IV-B Experimental Setup
IV-B1 Baselines
Since semantic segmentation can be regarded as a pixel-wise classification task, we compare our method with both classification and segmentation methods to provide a more comprehensive evaluation. The detailed baseline introduction is provided in the appendix.
IV-B2 Evaluation Metrics
Following [5], mean Intersection-over-Union ($mIoU$) is taken for evaluation:
$mIoU=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}},$ (12)
TABLE II: The table presents the 60-10 category configuration of AVSBench under the disjoint setting.
Disjoint Settings AVSBench-CI
IV-B3 Implementation Details of our methods
Our method builds upon the best-performing PLOP model combined with the memory. We have primarily conducted training and evaluation using ResNet-50 [40] pre-trained on ImageNet [41]. The ASPP module [9] is utilized as the fusion module. For input frames, we resize the resolution to $224\times 224$. The same data augmentation is applied as in [9], excluding memory data. The training batch size is set to 2 per GPU on 4 Nvidia L40 48GB GPUs. The training runs 30 epochs each task. For single-modal training, all steps are trained only using visual-modal data. For memory samples, 5 samples per class are selected for rehearsal. The memory dataset is shuffled together with the training dataset during training. The number of resampled samples is set to 20% of the total sample size. To be fair, all tasks share a common test set with all learned classes.
IV-B4 Implementation Details of baseline methods
For incremental classification methods: (1) Learning without forgetting (LWF) [35]: LWF distils the output differences between the old and current models. Our implementation of LwF follows [35]; distillation and cross-entropy losses share the same label space and classifier. (2) LwF multi-class (LWF-MC) [2]: LwF-MC utilizes multiple binary classifiers. Following the approach proposed in  [5], LWF-MC is implemented by combining two binary cross-entropy losses in a weighted manner. These losses are computed based on the ground truth labels and the probabilities predicted by the previous model $f_{\theta_{t-1}}$. (3) ILT: [36]: ILT employs a dual-space knowledge distillation strategy, including a distillation loss in the output space and an additional distillation loss in the feature space.
For incremental segmentation: (1) MiB [5]: MiB uses complete output space distillation and background uncertainty propagation. (2) PLOP [18]: PLOP proposes multi-scale pooling distillation to maintain spatial relationships at the feature level and uses entropy-based pseudo-labels to annotate background classes predicted by the old model. (3) EIR [22] is an instance rehearsal method for continual semantic segmentation, introduced in CVPR 2025, and represents the state-of-the-art (SOTA) in this field. In our work, We reproduced both the original EIR method and its PLOP-based variant, and adapted them to the continual audio-visual segmentation. Our experiments demonstrate that the PLOP-enhanced EIR outperforms the vanilla EIR approach. To ensure a fair comparison, we adopt the PLOP-based EIR method in our study.
Besides, the fine-tuning of AVSegFormer [37] is implemented based on ResNet-50. Additionally, fine-tuning each task as a baseline and offline training on all classes is provided as an upper bound for performance comparison.
IV-B5 The Details of Category
Tab. II present the 60-10 category learning sequence under the setting of disjoint in the AVSBench-CI dataset. For the setting of disjoint, we employ the Louvain algorithm to divide the 70-category dataset into bipartite and tripartite graphs. Classes with minimal overlapped are then allocated to distinct steps to form the disjoint dataset. The dataset was directly partitioned into steps based on sequential category order for the overlapped setting.
IV-C Main Results
@@
Tab. I illustrates the experiments of existing methods on AVSBench-CI. We use underlining to indicate the second-best performance. The upper bound represents the optimal performance when the model is directly trained on the target task. From left to right, task difficulty progressively increases, as more tasks lead to greater forgetting in the model. As reported in the results, our method achieves the best performance across all settings and demonstrates superior performance as the number of learning steps increases. On the more challenging 65-1 split, our method achieves signiffcantly better performance than traditional approaches. Despite incorporating audio, traditional continual semantic segmentation suffers significant forgetting due to its inability to effectively disentangle audio-visual interactions. Specifically, EIR exhibits consistently low performance. The primary reason is the poor rehearsal quality resulting from its inability to extract audio aligned with the synthesized content, which exacerbates modality entanglement and consequently leads to catastrophic forgetting. Thus, experimental results show that disentangling modalities is essential in audio-visual segmentation to mitigate catastrophic forgetting.
@@
IV-D Experiments on Transformer Architecture
To further validate the effectiveness of our method on Transformer-based architectures, we conduct additional experiments on the 60-10 and 60-5 settings using PVT (Pyramid Vision Transformer). The results in Tab. III demonstrate that our method continues to achieve competitive performance, even when applied to Transformer-based models, indicating its strong generalization capability across different architectural backbones.
TABLE III: The results of our method on the AVSBench-CI 60-10 task based on PVT
60-10
Disjoint Overlapped
Backbone 1-60 61-71 all 1-60 61-71 all
Ours (ResNet) 29.5 15.8 27.6 28.5 13.5 26.3
Ours (PVT) 33.7 34.7 33.9 35.1 15.6 32.4
IV-E Ablation Study
IV-E1 Effectiveness of MSS and CSR
@@
TABLE IV: Ablation Study on effectiveness of MSS and CSR.
@@
IV-E2 Number of rehearsal samples in MSS
Tab. V reports the results of the ablation study on the number of rehearsal samples per class. The results show that as the number of rehearsal samples increases, the forgetting of old classes gradually decreases. However, an excessive number of rehearsal samples can inhibit learning new samples. Therefore, we select five samples per class for rehearsal.
TABLE V: Ablation Study on the Number of rehearsal samples in MSS. We select 3, 5, and 7 samples per class using MSS.
@@ IV-E2 Number of rehearsal samples in MSS
IV-F Qualitative Analysis
IV-F1 Qualitive Analysis of AVSBench-CI
@@
Figure 7: Number of collision pairs. Highly colliding categories typically correspond to objects that co-occur. The categories with the highest collision rate are ”guitar” and ”man,” which aligns with real-world observations.
@@
IV-F2 Qualitive Analysis of AVSBench-CIM
Fig. 9 demonstrates a comparison between our method and previous methods on AVSBench-CIM, highlighting the superior performance of our method in scenarios requiring the segmentation of multiple targets. The figure presents three multi-target cases. In the first case, where the goal is to segment ”ukulele” and ”man,” our method achieves complete segmentation of both objects compared to previous methods while exhibiting significantly less class confusion. In the third case, while previous methods fail to segment the target object entirely, our method successfully segments most of the ”piano.” These examples further prove the superiority of our method in multi-target audio-visual segmentation tasks.
IV-F3 Qualitive Analysis of collision classes
IV-F4 Qualitive examples of Multi-modal semantic drift
To better understand the Multi-modal semantic drift task, we present two examples from the AVSBench-CI 60-10 task. The classes ”guitar” and ”drum” were learned in step 0, while ”violin” and ”woman” are to be learned in step 1. During the learning process of step 1, ”guitar” and ”drum” are labeled as background. This causes their corresponding audio to be associated with background semantics, leading to the multi-modal semantic drift.
Figure 10: Comparison of sample selection strategy. The image visualizes our method alongside the sample selection strategy based on maximum modality discrepancy, where samples selected exhibit greater consistency.
IV-F5 Effect analysis of Multi-modal Selection (MSS)
As shown in Fig. 10, the samples selected by MSS exhibit the following characteristics: (l) Unlike samples with multiple targets. MSS tends to favour samples with single targets. (2) MSS prefers samples where the target is consistently present. (3) MSS prioritizes samples with better alignment between thetarget audio and visual modalities. This phenomenon aligns with our initial hypothesis, as these three types of samples typically exhibit less multi-modal semantic drift, thereby aiding the model in better retaining knowledge of old classes.
Figure 11: Example of co-occurence. The top part of the figure shows that during training, the violin and bassoon frequently co-occur. As a result, during inference, the model mistakenly segments the bassoon as a violin.
IV-F6 Qualitive examples of Co-occurence confusion
In Figure 11, we present a example illustrating co-occurrence patterns in the data. During training, the classes ”violin” and ”bassoon” frequently co-occur across samples. At inference, the model correctly segments the spatial extent of the bassoon instance but erroneously assigns it the semantic label ”violin.” This observation suggests that while the model has effectively captured discriminative visual features, it exhibits semantic confusion when aligning visual inputs with their corresponding audio-derived class labels.
V Conclusion

### v2

@@ 4.2.1 Baselines
LwF Rebuffi et al. (2017) 10.1 25.1 12.3 7.1 19.0 8.8 1.5 9.7 2.6 1.5 12.6 3.0 1.3 0.7 1.3 1.3 4.5 1.6
LwF-MC Li and Hoiem (2018) 2.0 2.2 2.0 16.4 1.1 14.3 2.8 0.03 2.4 5.8 0.6 5.0 1.6 0.0 1.5 1.3 1.7 1.4
ILT Michieli and Zanuttigh (2019) 12.3 19.7 13.4 14.5 13.8 14.4 8.6 7.2 8.4 2.0 11.4 3.4 1.3 0.6 1.2 1.3 3.7 1.5
MiB Cermelli et al. (2020) 17.4 23.0 18.2 17.5 16.6 17.4 4.1 11.5 5.1 5.7 7.3 5.9 1.6 2.8 1.7 1.3 4.9 1.5
PLOP Douillard et al. (2021) 21.2 13.5 20.1 19.0 11.3 17.9 1.3 11.7 10.0 8.3 9.3 8.4 1.3 0.2 1.2 1.2 4.1 1.4
AVSegFormer Gao et al. (2024) 1.5 34.6 6.1 1.5 22.7 4.5 1.4 34.9 4.0 1.5 9.1 2.5 1.3 0.3 1.3 1.3 3.7 1.5
EIR Yin et al. (2025) 14.6 1.3 12.8 12.4 0.1 10.7 6.8 1.1 6.0 5.5 0.2 4.8 0.5 0.08 0.4 0.5 0.02 0.4
@@
4.3 Main Results
@@
Tab. 1 illustrates the experiments of existing methods on AVSBench-CI. We use underlining to indicate the second-best performance. The upper bound represents the optimal performance when the model is directly trained on the target task. From left to right, task difficulty progressively increases, as more tasks lead to greater forgetting in the model. As reported in the results, our method achieves the best performance across all settings and demonstrates superior performance as the number of learning steps increases. On the more challenging 65-1 split, our method achieves signiffcantly better performance than traditional approaches. Despite incorporating audio, traditional continual semantic segmentation suffers significant forgetting due to its inability to effectively disentangle audio-visual interactions. Specifically, EIR exhibits consistently low performance. The primary reason is the poor rehearsal quality resulting from its inability to extract audio aligned with the synthesized content, which exacerbates modality entanglement and consequently leads to catastrophic forgetting. Thus, experimental results show that disentangling modalities is essential in audio-visual segmentation to mitigate catastrophic forgetting.
@@
4.4 Ablation Study
4.4.1 Effectiveness of MSS and CSR
@@
Table 2: Ablation Study on effectiveness of MSS and CSR.
@@
Table 3: Ablation Study on the number of rehearsal samples in MSS. We select 3, 5, and 7 samples per class using MSS.
@@
4.5 Additional Ablation: Number of Rehearsal Samples
Tab. 3 reports the results of the ablation study on the number of rehearsal samples per class. The results show that as the number of rehearsal samples increases, the forgetting of old classes gradually decreases. However, an excessive number of rehearsal samples can inhibit learning new samples. Therefore, we select five samples per class for rehearsal.
4.6 Experiments on Transformer Architecture
To further validate the effectiveness of our method on Transformer-based architectures, we conduct additional experiments on the 60-10 and 60-5 settings using PVT (Pyramid Vision Transformer). The results in Tab. 5 demonstrate that our method continues to achieve competitive performance, even when applied to Transformer-based models, indicating its strong generalization capability across different architectural backbones.
Table 4: The results of our method on the AVSBench-CI 60-10 task based on PVT
60-10
Disjoint Overlapped
Backbone 1-60 61-71 all 1-60 61-71 all
Ours (ResNet) 29.5 15.8 27.6 28.5 13.5 26.3
Ours (PVT) 33.7 34.7 33.9 35.1 15.6 32.4
Table 5: The results of our method on the AVSBench-CI 60-5 task based on PVT.
60-5
Disjoint Overlapped
Backbone 1-60 61-71 all 1-60 61-71 all
Ours (ResNet) 26.2 11.6 24.2 24.3 10.4 22.4
Ours (PVT) 38.3 16.2 35.2 24.1 10.5 22.2
Moreover, results in Tab. 5 show that even with more powerful architectures, catastrophic forgetting still occurs under the 60-5 overlapped setting, highlighting the persistent challenge of knowledge retention in continual learning scenarios despite architectural advancements.
4.6.1 Qualitive Analysis of AVSBench-CI
5 Conclusion
Appendix A Appendix
A.1 Multimodal Learning
Multimodal learning Wei et al. (2024b); Wei and Hu (2024); Xiu et al. (2025); Guo et al. (2024) focuses on integrating information across diverse modalities and investigating the intricate interrelationships between them in various contexts. Wei Wei et al. (2024b) first estimates each modality’s learning status based on separability in its unimodal representation space, then uses this to softly initialize the corresponding unimodal encoder. MMPareto Wei and Hu (2024) employs gradient-based optimization to mitigate model bias towards specific modalities during training, thereby enhancing multimodal learning performance. To compare with more recent work, Finger Xiu et al. (2025) focuses on distinguishing foreground from background and transferring unimodal knowledge, while we focus on selecting consistent samples through modal contribution and replaying them according to collision frequency. From task level, Finger aims to seamlessly integrate new classes with limited incremental samples, while we focus on avoiding interference with old task knowledge when training on new tasks. Meanwhile, in contrast to Open-set AVS, continual learning AVS deals with learning from a continuous data stream under memory constraints, without revisiting past data.
A.2 Evaluation Metrics
Following Cermelli et al. (2020), mean Intersection-over-Union ($mIoU$) is taken for evaluation:
$mIoU=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}},$ (10)
A.3 Implementation Details
Our method builds upon the best-performing PLOP model combined with the memory. We have primarily conducted training and evaluation using ResNet-50 He et al. (2016) pre-trained on ImageNet Deng et al. (2009). The ASPP module Zhou et al. (2022) is utilized as the fusion module. For input frames, we resize the resolution to $224\times 224$. The same data augmentation is applied as in Zhou et al. (2022), excluding memory data. The training batch size is set to 2 per GPU on 4 Nvidia L40 48GB GPUs, and the training runs 30 epochs per task. For single-modal training, all steps are trained only using visual-modal data. For memory samples, 5 samples per class are selected for rehearsal, and the number of resampled samples is set to 20% of the total sample size. To be fair, all tasks share a common test set with all learned classes.
A.4 Baseline Methods
For incremental classification methods: (1) Learning without forgetting (LWF) Li and Hoiem (2018): LWF distils the output differences between the old and current models. Our implementation of LwF follows Li and Hoiem (2018); distillation and cross-entropy losses share the same label space and classifier. (2) LwF multi-class (LWF-MC) Rebuffi et al. (2017): LwF-MC utilizes multiple binary classifiers. Following the approach proposed in Cermelli et al. (2020), LWF-MC is implemented by combining two binary cross-entropy losses in a weighted manner. These losses are computed based on the ground truth labels and the probabilities predicted by the previous model $f_{\theta_{t-1}}$. (3) ILT Michieli and Zanuttigh (2019): ILT employs a dual-space knowledge distillation strategy, including a distillation loss in the output space and an additional distillation loss in the feature space.
For incremental segmentation: (1) MiB Cermelli et al. (2020): MiB uses complete output space distillation and background uncertainty propagation. (2) PLOP Douillard et al. (2021): PLOP proposes multi-scale pooling distillation to maintain spatial relationships at the feature level and uses entropy-based pseudo-labels to annotate background classes predicted by the old model. (3) EIR Yin et al. (2025) is an instance rehearsal method for continual semantic segmentation, introduced in CVPR 2025, and represents the state-of-the-art (SOTA) in this field. In our work, we reproduced both the original EIR method and its PLOP-based variant, and adapted them to the continual audio-visual segmentation. Our experiments demonstrate that the PLOP-enhanced EIR outperforms the vanilla EIR approach. To ensure a fair comparison, we adopt the PLOP-based EIR method in our study. Besides, the fine-tuning of AVSegFormer Gao et al. (2024) is implemented based on ResNet-50. Additionally, fine-tuning each task as a baseline and offline training on all classes is provided as an upper bound for performance comparison.
A.5 Additional Qualitative Results
A.5.1 Qualitative Analysis of AVSBench-CIM
Fig. 7 demonstrates a comparison between our method and previous methods on AVSBench-CIM, highlighting the superior performance of our method in scenarios requiring the segmentation of multiple targets. The figure presents three multi-target cases. In the first case, where the goal is to segment “ukulele” and “man,” our method achieves complete segmentation of both objects compared to previous methods while exhibiting significantly less class confusion. In the third case, while previous methods fail to segment the target object entirely, our method successfully segments most of the “piano.” These examples further prove the superiority of our method in multi-target audio-visual segmentation tasks.
A.5.2 The Details of Category
Tab. 7 present the 60-10 category learning sequence under the setting of disjoint in the AVSBench-CI dataset. For the setting of disjoint, we employ the Louvain algorithm to divide the 70-category dataset into bipartite and tripartite graphs. Classes with minimal overlapped are then allocated to distinct steps to form the disjoint dataset. The dataset was directly partitioned into steps based on sequential category order for the overlapped setting.
A.5.3 Qualitative Analysis of Collision Classes
A.5.4 Qualitative Analysis of Collision Classes
Table 6: The table presents the 60-10 category configuration of
AVSBench under the disjoint setting.
Disjoint Settings AVSBench-CI
@@ A.5.4 Qualitative Analysis of Collision Classes
Table 7: Number of collision pairs. Highly colliding categories typically correspond to objects that co-occur. The categories with the highest collision rate are “guitar” and “man,” which aligns with real-world observations.
A.5.5 Examples of Multi-modal Semantic Drift
@@
To better understand the Multi-modal semantic drift task, we present two examples from the AVSBench-CI 60-10 task. The classes “guitar” and “drum” were learned in step 0, while “violin” and “woman” are to be learned in step 1. During the learning process of step 1, “guitar” and “drum” are labeled as background. This causes their corresponding audio to be associated with background semantics, leading to the multi-modal semantic drift.
A.5.6 Effect Analysis of Multi-modal Sample Selection (MSS)
Figure 9: Comparison of sample selection strategy. The image visualizes our method alongside the sample selection strategy based on maximum modality discrepancy, where samples selected by MSS exhibit greater consistency.
As shown in Fig. 9, the samples selected by MSS exhibit the following characteristics: (1) Unlike samples with multiple targets, MSS tends to favour samples with single targets. (2) MSS prefers samples where the target is consistently present. (3) MSS prioritizes samples with better alignment between the target audio and visual modalities. This phenomenon aligns with our initial hypothesis, as these three types of samples typically exhibit less multi-modal semantic drift, thereby aiding the model in better retaining knowledge of old classes.